OpenStudy (anonymous):

What comes after 9/4 in a sequence?

OpenStudy (asnaseer):

you need more information. are you given the values before 9/4 in the sequence?

OpenStudy (anonymous):

Find the given term of each arithmetic sequence. 1/4,3/4,5/4,7/4,...;15th term

OpenStudy (asnaseer):

for example, in the sequence: 2,4,6,8 the value that comes after 8 would be 10 since we are adding 2 each time. whereas in the sequence: 0,4,8 the value that comes after 8 would be 12 as we are adding 4 each time. so you cannot answer "what comes after 8 in a sequence" with this other information.

OpenStudy (anonymous):

Find the given term of each arithmetic sequence. 1/4,3/4,5/4,7/4,...;15th term

OpenStudy (asnaseer):

ok, do you know what an "arithmetic sequence" is?

OpenStudy (anonymous):

Yes it is the difference between 2 numbers

OpenStudy (asnaseer):

sort of: it is a sequence where you add (or subtract) the same number from each term in a sequence. so for your problem, the first thing you should do is work out what the difference between any two consecutive terms is. what do you this is for your sequence?

OpenStudy (asnaseer):

what do you think this is for your sequence?

OpenStudy (anonymous):

It is 2 right?

OpenStudy (asnaseer):

again - almost, remember that each term is divided by 4, so the difference between each term is actually 2/4 (or 1/2).

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so after 7/4 would come 9/4

OpenStudy (asnaseer):

you can then use the standard formulae to find the nth term of an arithmetic sequence. do you know the formulae?

OpenStudy (asnaseer):

yes 9/4 would be after 7/4

OpenStudy (anonymous):

and after 9/4 would be 11/4 but im not sure how high it can go because i need to go to the 15 term, so i have to do it 15 times

OpenStudy (asnaseer):

you need to use the formulae for an arithmetic progression. are you aware of this formulae?

OpenStudy (asnaseer):

have you seen this formulae before?:\[a_n=a_1+(n-1)d\]

OpenStudy (anonymous):

and after 9/4 would be 11/4 but im not sure how high it can go because i need to go to the 15 term, so i have to do it 15 times

OpenStudy (anonymous):

o yes i have seen that formula but i forgot how to use it

OpenStudy (asnaseer):

ok, in this formulae, the first term of the sequence is a1. "d" is the difference between each term. "n" indicates which term you require. so, for your example, we would have:\[a_1=1/4\]\[d=1/2=0.5\]\[n=15\]so:\[a_15=1/4+(15-1)0.5\]

OpenStudy (anonymous):

oh ok

OpenStudy (asnaseer):

sorry:\[a_{15}=1/4+(15-1)/2\]

OpenStudy (anonymous):

yeah i understood but i have to go for a little while so thank you for teaching me the formula

OpenStudy (asnaseer):

no problem - glad to help out

OpenStudy (amistre64):

or simply, the 15th odd term divided by 4 :) (2(k)-1)/4 given that k=15